# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ 1 Introduction
Autoregressive Modelling has not matched diffusion models in image generation quality despite demonstrating strong performance in text [1]. The Visual Autoregressive Model (VAR) [2] bridges this gap by shifting from next-token to next-scale prediction, iteratively refining images by invoking an autoregressive transformer to predict each subsequent resolution. VAR is the first model to reach diffusion models in image quality while also improving runtime.
A key goal in autoregressive modeling is improving computational efficiency without sacrificing performance. Dynamic methods address this by allocating computation based on token complexity [3]. Currently, VAR still makes multiple static transformer calls per scale, ignoring varying complexity across tokens and scales. Exploiting these variations could enhance efficiency by focusing computation on complex regions and progressively reducing resources at finer scales [4]. Dynamic allocation at the scale level remains unexplored, presenting an opportunity to improve VAR’s efficiency. As a result, we adapt a dynamic Mixture-of-Experts (MoE) framework that routes experts based on token complexity, specifically for VAR’s coarse-to-fine structure [5]. We make the following contributions:
@@
    An empirical validation on ImageNet [6] showing 19% fewer FLOPs and 11% faster wall-clock time while staying within 1% Fréchet Inception Distance score [7] of the VAR depth 16 baseline.
@@ 2.1 Visual Autoregressive Modelling (VAR)
Rather than predicting tokens sequentially, VAR [8] redefines image generation as a coarse-to-fine process across multiple resolutions. An image is represented as a hierarchy of discrete token maps, ranging from a low to high resolution. The model then learns to autoregressively predict each finer scale conditioned on the coarser ones, progressively refining structure and texture until the full image is reconstructed. While this hierarchical design improves efficiency compared to raster-scan autoregression, related works for accelerating inference in VAR is discussed in Appendix A.
@@ 2.2 Dense-to-Dynamic-$k$ Mixture-of-Experts (D2DMoE)
From dense FFN to experts. In each Transformer block the two-layer FFN computes the hidden vector $\mathbf{h}=\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})$ and outputs $F(\mathbf{x})=\mathbf{W}_{2}\mathbf{h}+\mathbf{b}_{2}$. To make the activations $\mathbf{h}$ even sparser than with $\sigma$ set to ReLU, D2DMoE adds a Hoyer penalty [9] to the cross entropy loss:
@@
where $\mathbf{h}^{(\ell)}$ is the post-activation vector of FFN block $\ell$, $L$ is the number of blocks, and $\alpha$ controls sparsity strength. After sparsity-aware training, columns of $\mathbf{W}_{1}$ respond to subsets of tokens. By grouping these columns via $k$-means clustering into balanced clusters [10], D2DMoE rearranges the original FFN parameters into multiple distinct experts creating a Mixture-of-Experts layer.
@@ 3 Dynamic Mixture-of-Experts for Visual Autoregressive Model
Expert Construction. As detailed in section 2.2, at every Transformer layer, we replace the standard VAR’s FFN with a MoE block. Experts are built offline by clustering the sparse FFN weights from the ReLUfied model [11] finetuned with the Hoyer-norm regularisation (1). We then apply the same equation as in (2) to train a router that regresses the $\ell_{2}$-norm of each expert.
@@ 4 Results
Implementation Details. All experiments use the pre-trained VAR-d16 baseline, fine-tuned for two epochs with the sparse cross-entropy loss in Eq. (1). The full set of hyperparameters is provided in Appendix B. All experiments are conducted on NVIDIA A100 GPUs, measuring floating-point operations (FLOPs) using fvcore [^1] https://github.com/facebookresearch/fvcore and Fréchet Inception Distance (FID) using torch-fidelity [^2] https://github.com/toshas/torch-fidelity, with evaluation on ImageNet using 10,000 samples [6].
@@
$\Large\tau$-selection In Figure 3, “DMoE-VAR" employs a fixed $\tau$ for the last three scales. Varying $\tau$ produces a FID to FLOP trade-off curve, with the red cross denoting the configuration used in Figure 2. This thresholding is applied post-training, enabling flexible compute allocation at inference. Applying a uniform threshold to all scales ($\tau$ all scales”) performs worse in image quality and efficiency as errors from early scales propagate and low token counts fail to amortize the expert-loading overhead. As a result, the main gains come from sparsifying the later scales. To validate that token-wise routing is still required, we introduce “Pruned-FFN", which prunes low-activation weights in the last three scales. The sharp performance drop confirms the necessity of our dynamic thresholding.
@@ Appendix A Efficient Inference in VAR
To improve VAR’s efficiency, ScaleKV and Collaborative Decoding (CoDe) split the scale architecture into a “drafter" and a “refiner" stage. The drafter is responsible for generating a coarse, low-resolution version of the output sequence, while the refiner incrementally improves this draft to achieve high-quality results. ScaleKV reduces memory overhead by mapping transformer layers to these roles and compressing the key-value caches at higher resolutions [12]. CoDe similarly adopts a two-stage decoding structure, using a large-capacity drafter for coarse representations and a lightweight refiner for fine adjustments, enabling more targeted and efficient computation [13]. These architectures show that dividing decoding into specialized components can yield substantial efficiency gains without compromising output quality [14, 15].

### v2

# Jort Vincenti1, Metod Jazbec1, Guoxuan Xia2

@@ 1 Introduction
Autoregressive Modelling has not matched diffusion models in image generation quality despite demonstrating strong performance in text Ma et al. (2025). The Visual Autoregressive Model (VAR) Tian et al. (2024a) bridges this gap by shifting from next-token to next-scale prediction, iteratively refining images by invoking an autoregressive transformer to predict each subsequent resolution. VAR is the first model to reach diffusion models in image quality while also improving runtime.
A key goal in autoregressive modeling is improving computational efficiency without sacrificing performance. Dynamic methods address this by allocating computation based on token complexity Graves (2016). Currently, VAR still makes multiple static transformer calls per scale, ignoring varying complexity across tokens and scales. Exploiting these variations could enhance efficiency by focusing computation on complex regions and progressively reducing resources at finer scales Rao et al. (2021). Dynamic allocation at the scale level remains unexplored, presenting an opportunity to improve VAR’s efficiency. As a result, we adapt a dynamic Mixture-of-Experts (MoE) framework that routes experts based on token complexity, specifically for VAR’s coarse-to-fine structure Szatkowski et al. (2024). We make the following contributions:
@@
    An empirical validation on ImageNet Deng et al. (2009) showing 19% fewer FLOPs and 11% faster wall-clock time while staying within 1% Fréchet Inception Distance score Heusel et al. (2018) of the VAR depth 16 baseline.
@@ 2.1 Visual Autoregressive Modelling (VAR)
Rather than predicting tokens sequentially, VAR Tian et al. (2024b) redefines image generation as a coarse-to-fine process across multiple resolutions. An image is represented as a hierarchy of discrete token maps, ranging from a low to high resolution. The model then learns to autoregressively predict each finer scale conditioned on the coarser ones, progressively refining structure and texture until the full image is reconstructed. While this hierarchical design improves efficiency compared to raster-scan autoregression, related works for accelerating inference in VAR is discussed in Appendix A.
@@ 2.2 Dense-to-Dynamic-$k$ Mixture-of-Experts (D2DMoE)
From dense FFN to experts. In each Transformer block the two-layer FFN computes the hidden vector $\mathbf{h}=\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})$ and outputs $F(\mathbf{x})=\mathbf{W}_{2}\mathbf{h}+\mathbf{b}_{2}$. To make the activations $\mathbf{h}$ even sparser than with $\sigma$ set to ReLU, D2DMoE adds a Hoyer penalty Hoyer (2004) to the cross entropy loss:
@@
where $\mathbf{h}^{(\ell)}$ is the post-activation vector of FFN block $\ell$, $L$ is the number of blocks, and $\alpha$ controls sparsity strength. After sparsity-aware training, columns of $\mathbf{W}_{1}$ respond to subsets of tokens. By grouping these columns via $k$-means clustering into balanced clusters Zhang et al. (2022), D2DMoE rearranges the original FFN parameters into multiple distinct experts creating a Mixture-of-Experts layer.
@@ 3 Dynamic Mixture-of-Experts for Visual Autoregressive Model
Expert Construction. As detailed in section 2.2, at every Transformer layer, we replace the standard VAR’s FFN with a MoE block. Experts are built offline by clustering the sparse FFN weights from the ReLUfied model Mirzadeh et al. (2023) finetuned with the Hoyer-norm regularisation (1). We then apply the same equation as in (2) to train a router that regresses the $\ell_{2}$-norm of each expert.
@@ 4 Results
Implementation Details. All experiments use the pre-trained VAR-d16 baseline, fine-tuned for two epochs with the sparse cross-entropy loss in Eq. (1). The full set of hyperparameters is provided in Appendix B. All experiments are conducted on NVIDIA A100 GPUs, measuring floating-point operations (FLOPs) using fvcore [^1] https://github.com/facebookresearch/fvcore and Fréchet Inception Distance (FID) using torch-fidelity [^2] https://github.com/toshas/torch-fidelity, with evaluation on ImageNet using 10,000 samples Deng et al. (2009).
@@
$\tau$-selection In Figure 3, “DMoE-VAR" employs a fixed $\tau$ for the last three scales. Varying $\tau$ produces a FID to FLOP trade-off curve, with the red cross denoting the configuration used in Figure 2. This thresholding is applied post-training, enabling flexible compute allocation at inference. Applying a uniform threshold to all scales ($\tau$ all scales”) performs worse in image quality and efficiency as errors from early scales propagate and low token counts fail to amortize the expert-loading overhead. As a result, the main gains come from sparsifying the later scales. To validate that token-wise routing is still required, we introduce “Pruned-FFN", which prunes low-activation weights in the last three scales. The sharp performance drop confirms the necessity of our dynamic thresholding.
@@ Appendix A Efficient Inference in VAR
To improve VAR’s efficiency, ScaleKV and Collaborative Decoding (CoDe) split the scale architecture into a “drafter" and a “refiner" stage. The drafter is responsible for generating a coarse, low-resolution version of the output sequence, while the refiner incrementally improves this draft to achieve high-quality results. ScaleKV reduces memory overhead by mapping transformer layers to these roles and compressing the key-value caches at higher resolutions Li et al. (2025). CoDe similarly adopts a two-stage decoding structure, using a large-capacity drafter for coarse representations and a lightweight refiner for fine adjustments, enabling more targeted and efficient computation Chen et al. (2024). These architectures show that dividing decoding into specialized components can yield substantial efficiency gains without compromising output quality Xie et al. (2024), Gong et al. (2025).